### v1

@@ 1 Introduction
Large Language Models (LLMs) have triggered a paradigm shift in AI, excelling at tasks such as instruction following, multi-turn dialogue, and code generation openai2023gpt4. Most current LLMs are autoregressive, decoder-only Transformers trained to predict the next token under a causal attention mask dubey2024llama; zhang2022opt. While training is parallelizable via teacher forcing, inference inherits a rigid sequential dependency, becoming a key efficiency bottleneck. Moreover, such models cannot exploit bidirectional context, which limits performance on tasks such as text infilling, editing, and global planning hu2024bai; song2025seeddiffusion.
@@
As a promising alternative, diffusion language models (DLMs) have drawn attention for their potential in text generation. Inspired by the success of diffusion models in image, audio, and video synthesis ho2020ddpm, DLMs iteratively denoise random or masked token sequences into coherent text gong2023diffuseq; li2022diffusionlm. This process updates all token positions in parallel per step, removing the token-level sequential dependency inherent to autoregressive decoding. Closed-source models, primarily in code generation, report up to $10\times$ higher throughput than autoregressive models while preserving quality deepmind2025geminidiffusion; khanna2025mercury; song2025seeddiffusion. However, open-source DLMs remain significantly slower nie2025large; ye2025dream7b, largely because (1) bidirectional attention prohibits standard KV caching and (2) they demand many refinement steps, often proportional to sequence length, to reach high fidelity kim2025beyond.
@@
Many recent works aim to accelerate DLM inference. One primary strategy leverages block-wise decoding with KV caching, either via approximate caching with periodic updates hu2025freecache; liu2025dllmcache; ma2025dkvcache; wu2025fastdllm or by fine-tuning the model into a fully block-wise causal architecture so that caching becomes natural wang2025d2f; wu2025fastdllmv2. The other major line of work focuses on parallel sampling algorithms, which effectively decode multiple tokens per denoising step and reduce the total number of iterations benhamu2025ebsampler; wu2025fastdllm.
@@
Turning to diffusion-based applications in vision domains, many works have successfully reduced denoising to few-step or even one-step generation yin2024dmd; yin2025causvid. A prominent technique in that realm is consistency modeling: consistency models are trained to minimize discrepancies between adjacent denoising states along the probability flow ODE, enabling near one-step generation song2023consistency; song2023improved. Although originally developed for continuous data, the consistency paradigm has begun to influence discrete settings, including parallel decoding for autoregressive language models kou2024cllms and masked diffusion frameworks xu2025unicms, indicating its strong promise in discrete language domains.
@@
To drive DLMs to converge to coherent text in far fewer refinement steps, we present CDLM, a training-based acceleration method that integrates Consistency modeling into Diffusion Language Models. Concretely, the student is trained to finalize multiple tokens per step by distilling from a fully bidirectional teacher and by aligning its predictions between a less-informed state and its block-completion state along the decoding trajectory. We impose a block-wise causal attention mask and jointly optimize three objectives. At inference time, the model decodes sequentially over blocks while applying confidence-thresholded parallel decoding within each block. This design directly tackles two core bottlenecks of current DLMs: (i) excessive sampling steps (mitigated by distillation and consistency) and (ii) cache inefficiency (removed by block-wise causality). In particular, we make the following contributions:
@@
  • Consistency-guided distillation. We bring consistency modeling to DLMs by defining a token-level decoding trajectory and coupling it with distillation from a fully bidirectional teacher. This joint objective provides explicit supervision for multi-token finalization while keeping predictions stable within each block (see Sections 3 and 4.2).
@@
  • Cache-friendly block causality. By fine-tuning with a block-wise causal attention mask (Figure 2), we enable bidirectional DLMs to support block-wise KV caching and early stopping at block boundaries (see Sections 4.1 and 4.3).
@@
  • End-to-end speedups. With 8 hours of training for Dream and 16 hours for LLaDA, CDLM cuts refinement steps by $3.4{\times}$$7.9{\times}$ and latency by $3.6{\times}$$14.5{\times}$ while maintaining accuracy across math and coding benchmarks. It also surpasses equal-size autoregressive LLMs in tokens-per-second (see Section 5.2).
@@ 2.1 Diffusion Language Models
Diffusion models ho2020ddpm have been successfully applied across modalities such as text-conditioned image generation ramesh2022hierarchical, audio synthesis chen2021wavegrad, and video generation esser2023sceneguided. In language modeling, diffusion language models (DLMs) generate fixed-length sequences via iterative refinement: at each step, multiple token positions are updated in parallel and finalized gong2023diffuseq; li2022diffusionlm. Early approaches operated in the continuous embedding space gao2024embedding; gong2023diffuseq, but more recent models adopt masked diffusion models (MDMs) that work directly in the discrete token space gong2025scalingdiffusion; nie2025scalingmdm. They begin from a fully masked sequence and iteratively unmask tokens according to a scheduling policy. In this paper, we treat MDMs as the representative class of DLMs, especially Dream ye2025dream7b and LLaDA nie2025large. Architecturally, DLMs commonly employ Transformer backbones with bidirectional (non-causal) attention, permitting full-context interactions nie2025large. Despite their parallel update capability, DLMs are known to be slow for two main reasons. First, full bidirectional attention at every step renders them incompatible with standard KV caching. Second, high-quality generation typically demands a large number of refinement steps comparable to the sequence length, making inference inefficient kim2025beyond. In this work, we propose a fine-tuning-based solution that addresses both limitations.
@@ 2.2 Efficient Inference in Diffusion Language Models
Inference in DLMs is an active area of research with significant room for improvement. Among training-free acceleration methods, one line of work exploits approximate caching via block-wise decoding, refreshing static regions periodically ma2025dkvcache; wu2025fastdllm. Others speed up parallel sampling by applying confidence thresholds wu2025fastdllm or by leveraging lightweight autoregressive models as guidance hu2025freecache. In contrast, training-based approaches focus on improving the intrinsic efficiency of DLMs through architectural or objective-level modifications. Some recent works fine-tune models to introduce block-wise causality, enabling standard KV caching at the block level while retaining diffusion-style updates within each block wang2025d2f; wu2025fastdllmv2. Others design objectives that promote faster convergence by encouraging high-confidence token prediction chen2025dparallel. Our method is training-based: we induce block-wise causality and cut refinement steps via fine-tuning.
@@ 2.3 Knowledge Distillation in Language Models
Knowledge distillation (KD) hinton2015distilling transfers knowledge from a large, powerful teacher to a smaller, more efficient student. In early LLM work, KD was largely black-box, where students imitated textual outputs from closed-source APIs fu2023specializing. With the rise of open-source LLMs, white-box distillation became feasible, enabling direct supervision from the teacher’s internal representations, such as logits or embeddings agarwal2024onpolicy; gu2023knowledge. Beyond language modeling, distillation has also been successfully applied to diffusion-based image and video generation to cut sampling steps while preserving quality yin2024dmd; yin2025causvid. Our method can similarly be viewed as a form of self-distillation: we distill knowledge from a teacher to a student of identical size and model type, but with different architectural design (a fully bidirectional teacher vs. a block-wise causal student). We adopt a white-box setup with rich intermediate supervision, enabling efficient and stable knowledge transfer. Like deschenaux2025sdtt and xu2025dinfer, we reduce decoding steps by distilling multi-step diffusion trajectories into jumps between non-consecutive denoising states. Our novelty is to instantiate these jumps in a block-causal student, and to train it with consistency-oriented objectives.
@@ 2.4 Consistency Models and Language Models
Consistency models song2023consistency were introduced to overcome the slow, iterative sampling process of diffusion models. They learn a consistency property: any intermediate state along a probability flow ODE trajectory can be mapped directly back to its origin (the clean data), enabling near one-step generation. song2023improved. While originally developed for continuous domains such as vision, recent work has explored adapting this idea to discrete language modeling kou2024cllms; xu2025unicms. Rather than explicitly modeling the ODE solver or its continuous trajectory, these methods reinterpret the notion of mapping intermediate states to final solutions within discrete token trajectories. For instance, autoregressive language models, despite lacking an explicit diffusion trajectory, realize this principle via Jacobi-style parallel decoding, treating iterative drafts as intermediate states santilli2023accelerating. Our approach adopts this paradigm in DLMs, where trajectories arise naturally from iterative denoising. Conceptually, our objective can be viewed as an empirical generalization of the original consistency objective defined over continuous ODE trajectories, adapted to discrete, token-level diffusion processes.
@@ 3 Preliminary: Inference in Diffusion Language Models
Following prior formulations austin2021structured, the reverse-time transition $q_{s|t}(\mathbf{x}_{s}\mid\mathbf{x}_{t})$ is defined for $0\leq s<t\leq 1$ and factorizes across tokens. For each token $i$, the transition probability is defined as:
@@
where $q_{0|t}$ is the predictive distribution induced by $p_{\theta}(\mathbf{x}_{0}\mid\mathbf{x}_{t},c)$. This formulation captures three possibilities: (1) an already-unmasked token remains unchanged, (2) a masked token remains masked, or (3) a masked token is unmasked to a new token.
@@
To instantiate the probabilistic distribution $q_{s|t}$ into an actual generation procedure, practical deterministic strategies are employed. A common choice is low-confidence remasking, where instead of stochastic sampling, the model greedily selects the tokens to unmask nie2025large. Based on the confidence scores from $p_{\theta}(\mathbf{x}_{0}\mid\mathbf{x}_{t},c)$, the top-$m$ masked tokens with the highest probabilities are revealed, while all others remain masked. In practice, this process is often constrained to operate within blocks of tokens, called block-wise decoding, limiting unmasking to localized regions of the sequence. This prevents unnatural early generation (e.g., premature <endoftext> tokens) nie2025large and, more importantly, enables KV caching for finalized blocks.
@@ 4 Methodology
We present Consistency Diffusion Language Models (CDLM), a training scheme that accelerates DLM inference without sacrificing quality. We first describe how we collect teacher decoding trajectories and store compact hidden states for logit reconstruction, then introduce the distillation, consistency, and DLM objectives, and finally detail our confidence-thresholded block-wise inference.
@@
4.1 Trajectory Collection for Diffusion Language Models
Data Generation
@@ 4.1 Trajectory Collection for Diffusion Language Models
Prior work reports that block-wise decoding is effective for instruction-tuned models and attains high quality when the number of steps matches the generation length nie2025large. Accordingly, we adopt block-wise decoding with the number of sampling steps equal to the generation length ($N{=}L_{g}$) and finalize exactly one token per step within the current block (the highest-confidence token $y_{i}$). This configuration places the teacher at its most performant operating point and yields higher-quality trajectories. For white-box distillation, we store not only $\mathcal{T}_{x}$ but also the states required to reconstruct logits. Concretely, we record the last hidden states at the moments when tokens are finalized.
Data Augmentation
Effective training benefits from trajectories that are both diverse and reliable. Unlike autoregressive models, where temperature only influences token selection, in DLMs, the temperature also changes the order in which tokens are revealed gong2025diffucoder which allows further diversity in the generated data. We therefore augment each prompt with multiple trajectories generated at different temperatures. Additional details on dataset construction are provided in Appendix A.1.
@@ 4.2 Training
Here, $q_{\phi^{-}}$ denotes a stop-gradient target (detached from backpropagation) for stable training song2023consistency, and $D_{\mathrm{KL}}$ is the forward KL divergence between two distributions. Intuitively, Eq. (5) aligns the student at a less-informed state with itself at a more-informed state, but only over $\mathcal{S}_{y}$ where predictions are well-defined. Since DLMs are trained to predict only [MASK] tokens, restricting the loss to still-masked indices avoids ill-defined supervision and helps the model jump several steps in a stable, consistent manner.
@@
We include an auxiliary loss, identical to the masked denoising objective used in standard DLM pre-training nie2025large; ye2025dream7b, to preserve the model’s mask prediction capability. For a prompt $x$ and its ground-truth response $\hat{\mathbf{y}}$, we randomly sample a masking ratio $t\sim\mathcal{U}[0,1]$ and independently mask each token with a probability $t$ to obtain the masked sequence $\hat{\mathbf{y}}_{t}$. The resulting objective is
@@ 4.3 Inference
At inference time, the student decodes block-wise under a block-causal mask (Figure 2), reusing the KV cache for the prompt and all previously finalized blocks. Within each block, we apply confidence-thresholded parallel finalization: at each step, among the masked positions of the current block, we reveal every token whose confidence exceeds a threshold $\tau_{\text{conf}}$, following Fast-dLLM wu2025fastdllm. Larger $\tau_{\text{conf}}$ yields more conservative decoding (typically higher quality but slower convergence), whereas smaller $\tau_{\text{conf}}$ is more aggressive (faster convergence) at the cost of potential accuracy loss. We also adopt early stopping: decoding terminates once an <endoftext> token is produced within the current block, analogous to AR decoding. We intentionally avoid additional heuristics such as inter-block parallelism wang2025d2f, which introduce extra hyperparameters whose optimal values are task- and domain-dependent.

### v2

@@ 1 Introduction
Large Language Models (LLMs) have triggered a paradigm shift in AI, excelling at tasks such as instruction following, multi-turn dialogue, and code generation Achiam et al. (2023). Most current LLMs are autoregressive, decoder-only Transformers trained to predict the next token under a causal attention mask Grattafiori et al. (2024); Zhang et al. (2022). While training is parallelizable via teacher forcing, inference inherits a rigid sequential dependency, becoming a key efficiency bottleneck. Moreover, such models cannot exploit bidirectional context, which limits performance on tasks such as text infilling, editing, and global planning Hu et al. (2024); Song et al. (2025).
@@
As a promising alternative, diffusion language models (DLMs) have drawn attention for their potential in text generation. Inspired by the success of diffusion models in image, audio, and video synthesis Ho et al. (2020), DLMs iteratively denoise random or masked token sequences into coherent text Gong et al. (2023); Li et al. (2022). This process updates all token positions in parallel per step, removing the token-level sequential dependency inherent to autoregressive decoding. Closed-source models, primarily in code generation, report up to $10\times$ higher throughput than autoregressive models while preserving quality Google DeepMind (2025); Khanna et al. (2025); Song et al. (2025). However, open-source DLMs remain significantly slower Nie et al. (2025b); Ye et al. (2025), largely because (i) bidirectional attention prohibits standard KV caching and (ii) they demand many refinement steps, often proportional to sequence length, to reach high fidelity Kim et al. (2025).
@@
Many recent works aim to accelerate DLM inference. One primary strategy leverages block-wise decoding with KV caching, either via approximate caching with periodic updates Hu et al. (2025); Liu et al. (2025b); Ma et al. (2025a); Wu et al. (2025b) or by fine-tuning the model into a block-wise causal architecture so that caching becomes natural Wang et al. (2025); Wu et al. (2025a). The other major line of work focuses on parallel sampling algorithms, which effectively decode multiple tokens per denoising step and reduce the total number of iterations Ben-Hamu et al. (2025); Wu et al. (2025b).
@@
In diffusion models for vision domains, many works have successfully reduced denoising to few-step or even one-step generation Yin et al. (2024; 2025). A key technique behind this progress is consistency modeling, which trains models to produce consistent predictions across adjacent denoising states along the probability flow trajectory Song et al. (2023); Song and Dhariwal (2023). Although originally developed for continuous data, this paradigm has begun to influence discrete settings, including parallel decoding for autoregressive language models Kou et al. (2024) and masked diffusion frameworks Xu et al. (2025), highlighting its promise for language diffusion models.
@@
In this work, we present CDLM, a training-based acceleration method that integrates Consistency modeling Song et al. (2023); Song and Dhariwal (2023) into Diffusion Language Models. CDLM trains a block-causal student via distillation from a bidirectional teacher to finalize multiple tokens per refinement step, and applies confidence-thresholded parallel decoding within each block at inference time. This design directly addresses two core bottlenecks of current DLMs: (i) excessive refinement steps (mitigated by distillation and consistency) and (ii) cache inefficiency (removed by block-wise causality). In particular, we make the following contributions:
@@
  • Consistency-guided distillation. We bring consistency modeling to DLMs by defining a token-level decoding trajectory and coupling it with distillation from a fully bidirectional teacher. This joint objective provides explicit supervision for multi-token finalization while aligning predictions between a less-informed state and a block-completion state within each block (Sections 34.2).
@@
  • Cache-friendly block causality. By fine-tuning with a block-wise causal attention mask (Figure 2), we make bidirectional DLMs compatible with block-wise KV caching and early stopping at block boundaries ( Sections 4.14.3). We also provide a detailed system-level analysis of block-causal DLMs to characterize their efficiency and scalability (Section 5.4).
@@
  • End-to-end speedups. With 8 hours of training for Dream and 16 hours for LLaDA, CDLM reduces refinement steps by $3.4{\times}$$7.9{\times}$ and latency by $3.6{\times}$$14.5{\times}$ while maintaining accuracy across math and coding benchmarks. It also surpasses equal-size autoregressive LLMs in tokens-per-second (Section 5.2).
@@ 2.1 Diffusion Language Models
Diffusion models Ho et al. (2020) have been successfully applied across modalities such as text-conditioned image generation Ramesh et al. (2022), audio synthesis Chen et al. (2021b), and video generation Esser et al. (2023). Early Diffusion Language Models (DLMs) operated in the continuous embedding space Gao et al. (2024); Gong et al. (2023), but more recent models adopt masked diffusion models (MDMs) that work directly in the discrete token space Gong et al. (2025a); Nie et al. (2025a). These models begin from a fully masked sequence and iteratively unmask tokens according to a scheduling policy. Architecturally, DLMs commonly employ Transformer backbones with bidirectional (non-causal) attention, permitting full-context interactions Nie et al. (2025b). Despite their parallel update capability, DLMs are known to be slow for two main reasons. First, full bidirectional attention at every denoising step makes DLMs incompatible with standard KV caching. Second, high-quality generation demands a large number of denoising steps comparable to the sequence length, making inference inefficient Kim et al. (2025). In this work, we propose a fine-tuning-based solution that addresses both limitations.
@@ 2.2 Efficient Inference in Diffusion Language Models
Inference in DLMs is an active area of research with significant room for improvement. Among training-free acceleration methods, one line of work exploits approximate caching via block-wise decoding, only recomputing the active block while reusing stale key–value states for inactive blocks Ma et al. (2025a); Wu et al. (2025b). Others speed up parallel sampling by applying confidence thresholds Wu et al. (2025b) or by leveraging lightweight autoregressive models as guidance Hu et al. (2025), while recent analysis suggests that effective parallelism is difficult to adapt reliably at inference time Kang et al. (2025). In contrast, training-based approaches focus on improving the intrinsic efficiency of DLMs through architectural or objective-level modifications. Some recent works fine-tune pretrained AR models or DLMs to introduce block-wise causality, enabling standard KV caching at the block level while retaining diffusion-style updates within each block Wang et al. (2025); Wu et al. (2025a). Others design objectives that promote faster convergence by encouraging high-confidence token prediction Chen et al. (2025). Our method is training-based: we jointly induce block-wise causality and reduce refinement steps through consistency-guided fine-tuning.
@@ 2.3 Knowledge Distillation in Language Models
Knowledge distillation (KD) Hinton et al. (2015) transfers knowledge from a large, powerful teacher to a smaller, more efficient student. Early work on LLM distillation primarily relied on black-box KD, where students imitated textual outputs from closed-source APIs Fu et al. (2023). With the rise of open-source LLMs, white-box distillation has become prevalent, enabling richer supervision from the teacher’s internal representations, such as logits or embeddings Agarwal et al. (2024); Gu et al. (2023). Beyond language, distillation has also been applied to diffusion-based image and video generation to reduce sampling steps while preserving quality Yin et al. (2024; 2025). Our method can similarly be viewed as a form of self-distillation, where knowledge is transferred between models of identical size and type but with different architectural designs (a fully bidirectional teacher and a block-wise causal student). Like Deschenaux and Gulcehre (2025) and Ma et al. (2025b), we reduce decoding steps by distilling multi-step diffusion trajectories into jumps between non-consecutive denoising states. Our novelty lies in instantiating these jumps within a block-causal student and training it with consistency-oriented objectives.
@@ 2.4 Consistency Models and Language Models
Consistency models Song et al. (2023) were introduced to overcome the slow, iterative sampling process of diffusion models. They learn a consistency property: any intermediate state along a probability flow ODE trajectory can be mapped directly back to its origin (the clean data) Song and Dhariwal (2023). This is achieved by minimizing discrepancies between predictions at adjacent denoising states. While originally developed for continuous domains such as vision, recent work has explored adapting this idea to discrete language modeling Kou et al. (2024); Xu et al. (2025). Rather than explicitly modeling the ODE solver or its continuous trajectory, these methods reinterpret the notion of mapping intermediate states to final solutions within discrete token trajectories Santilli et al. (2023). Our approach adopts this paradigm in DLMs, where trajectories arise naturally from iterative denoising. Conceptually, our objective can be viewed as an empirical generalization of the original consistency objective defined over continuous ODE trajectories, adapted to discrete, token-level diffusion processes.
@@ 3 Preliminary: Inference in Diffusion Language Models
Following prior formulations Austin et al. (2021a), the reverse-time transition $q_{s|t}(\mathbf{x}_{s}\mid\mathbf{x}_{t})$ is defined for $0\leq s<t\leq 1$ and factorizes across tokens. For each token $i$, the transition probability is defined as:
@@
where $q_{0|t}$ is the predictive distribution induced by $p_{\theta}(\mathbf{x}_{0}\mid\mathbf{x}_{t},c)$. This formulation reflects three token-level transitions: preserving an unmasked token, keeping a token masked, or unmasking it to a new token.
@@
To instantiate the probabilistic distribution $q_{s|t}$ into an actual generation procedure, practical deterministic strategies are employed. A common choice is low-confidence remasking, where instead of stochastic sampling, the model greedily selects the tokens to unmask Nie et al. (2025b). Based on the confidence scores from $p_{\theta}(\mathbf{x}_{0}\mid\mathbf{x}_{t},c)$, the top-$m$ masked tokens with the highest probabilities are revealed, while all others remain masked. In practice, this process is often constrained to operate within blocks of tokens, called block-wise decoding, limiting unmasking to localized regions of the sequence. This prevents unnatural early generation Nie et al. (2025b) and, more importantly, enables KV caching for finalized blocks. Parallel denoising implicitly assumes independence among simultaneously finalized tokens, an idealized assumption that can degrade generation quality as more tokens are unmasked at once Liu et al. (2025a).
@@ 4 Methodology
We present Consistency Diffusion Language Models (CDLM), a training scheme that accelerates DLM inference without sacrificing quality. We describe trajectory and hidden-state collection, our three-objective training, and the inference strategy.
@@
4.1 Trajectory Collection for CDLM
@@
Prior work reports that block-wise decoding is effective for instruction-tuned models and attains high quality when the number of steps matches the generation length Nie et al. (2025b). Accordingly, we adopt block-wise decoding with the number of sampling steps equal to the generation length ($N{=}L_{g}$) and finalize exactly one token per step within the current block (the highest-confidence token $y_{i}$). This configuration places the teacher at its most performant operating point and yields higher-quality trajectories. For white-box distillation, we store not only $\mathcal{T}_{x}$ but also the states required to reconstruct logits. Concretely, we record the last hidden states at the moments when tokens are finalized. We also augment each prompt with multiple trajectories generated at different temperatures. Additional details on dataset construction are provided in Appendix A.1.
@@ 4.2 CDLM Training
Here, $q_{\phi^{-}}$ denotes a stop-gradient target (detached from backpropagation) for stable training Song et al. (2023), and $D_{\mathrm{KL}}$ is the forward KL divergence between two distributions. Intuitively, Eq. (5) aligns the student at a less-informed state with itself at a more-informed state, encouraging stable multi-step jumps along the decoding trajectory. Since DLMs are trained to predict only [MASK] tokens, restricting the loss to still-masked positions ($\mathcal{S}_{y}$) avoids ill-defined supervision.
@@
We include an auxiliary loss, identical to the masked denoising objective used in standard DLM pre-training Nie et al. (2025b); Ye et al. (2025), to preserve the model’s mask prediction capability. For a prompt $x$ and its ground-truth response $\hat{\mathbf{y}}$, we randomly sample a masking ratio $t\sim\mathcal{U}[0,1]$ and independently mask each token with a probability $t$ to obtain the masked sequence $\hat{\mathbf{y}}_{t}$. The resulting objective is
@@ 4.3 Inference
At inference time, the student decodes block-wise under a block-causal mask (Figure 2), reusing the KV cache for the prompt and all previously finalized blocks. Within each block, we apply confidence-thresholded parallel finalization: at each step, among the masked positions of the current block, we reveal every token whose confidence exceeds a threshold $\tau_{\text{conf}}$, following Fast-dLLM Wu et al. (2025b). We also adopt early stopping: decoding terminates once an <endoftext> token is produced within the current block, analogous to AR decoding. We intentionally avoid additional heuristics such as inter-block parallelism Wang et al. (2025), which introduce extra hyperparameters whose optimal values are task- and domain-dependent.